OpenStudy (anonymous):

how do you simplify the following expression? square root of -144 square root of -48

OpenStudy (anonymous):

\[x=\sqrt{-144}=+-12 i\]

OpenStudy (anonymous):

how sis you do that i dont understand

OpenStudy (anonymous):

i is iota used for defining imaginary numbers...as such no real numbers can define sqrt of a negative number

OpenStudy (anonymous):

i m not your sis...siisy

OpenStudy (anonymous):

sorry ment did

OpenStudy (markrijckenberg):

The s is right next to the d on an azerty keyboard. It must have been an accident indeed :-)

OpenStudy (anonymous):

can you explain it a little better I think that person got mad and I didnt mean to mae them mad

OpenStudy (anonymous):

oops make

OpenStudy (anonymous):

these two dont have anything between them. it is sqrt-144sqrt-48

OpenStudy (phi):

is it \[ \sqrt{-144}\sqrt{-48} \]

OpenStudy (anonymous):

yes

OpenStudy (markrijckenberg):

http://en.wikipedia.org/wiki/Square_root#Square_roots_of_negative_and_complex_numbers √-144=√-1 * √144 = +-i * 12 √-48 = √-1 * √48 = +-i * √48

OpenStudy (phi):

so you are multiplying two square roots. First notice that they are negative (that means you will get an imaginary number) write them as \[ \sqrt{144}\sqrt{-1}\sqrt{48}\sqrt{-1} \] the sqrt(-1) is called i ("i" for imaginary") you are multiplying, so you can change the order (example:2*3 is the same as 3*2) \[ \sqrt{144}\sqrt{48}\sqrt{-1}\sqrt{-1} \] now simplify: what is the sqrt(144) (it is a whole number)

OpenStudy (anonymous):

ok 144 is 12 and not sure what 48 is

OpenStudy (anonymous):

I dont think there is one for 48

OpenStudy (anonymous):

\[4\sqrt{3}\]

OpenStudy (phi):

to simplify 48, the easiest way (other than memorizing things) is factor it into its prime factors: first divide by 2 2 * 24 2 * 2*12 2*2*2*6 2*2*2*2*3 now look for pairs of numbers. here you have two pairs of 2's (2*2)*(2*2)*3 "pull out" the 2 pair from the square root sqrt(2*2 * 2*2) * sqrt(3) 2*2*sqrt(3) 4*sqrt(3) now do sqrt(-1)* sqrt(-1) use the rule that sqrt(x)*sqrt(x) = x

OpenStudy (anonymous):

i m not telling him to memorize if u knows abt iota then he must be knowing about factoization of a number too...

OpenStudy (anonymous):

yes i just learned that. didnt think that was involvd in this. sorry thought about if afterwards

OpenStudy (anonymous):

sorry I am writing this down so i can have an example of what to do

OpenStudy (phi):

now do sqrt(-1)* sqrt(-1) use the rule that sqrt(x)*sqrt(x) = x

OpenStudy (anonymous):

not sure what yor are saying sorry this is all new work for me.

OpenStudy (phi):

there is rule (or definition) that says sqrt(x) * sqrt(x) = x for example, sqrt(4)*sqrt(4)= 4 2 * 2 = 4 it worked sqrt(9)*sqrt(9)= 9 3 * 3 = 9 yes it works, sqrt(-1)*sqrt(-1)= ?

OpenStudy (anonymous):

=1

OpenStudy (phi):

look at the rule more carefully

OpenStudy (anonymous):

I thought that two negatives = a positive. so -1*-1 = 1 is that what you are looking for

OpenStudy (phi):

yes, -1* -1= +1 but we are doing sqrt(-1)*sqrt(-1) and we have to use the rule sqrt(x) * sqrt(x) = x

OpenStudy (anonymous):

or are you looking for 12*4 sqrt 3 I think I am not understanding what you are saying

OpenStudy (anonymous):

sorry i am not one to just get an answer i want to know how to do it

OpenStudy (phi):

OK, we can go over the why in a minute, but first, let's get the answer to your question we have \[ 12\cdot 4\sqrt{3}\cdot \sqrt{-1}\cdot\sqrt{-1} \] we can simplify the last part by using sqrt(x)*sqrt(x) = x so what is \[ \sqrt{-1}\cdot\sqrt{-1} = ?\]

OpenStudy (anonymous):

ok I am writing it down.

OpenStudy (anonymous):

um I am guessing would it be i for the imaginary number

OpenStudy (phi):

no, i is short hand for \( \sqrt{-1} \) to save typing the square root of -1 all the time. there is only one way to know what \[ \sqrt{-1}\cdot\sqrt{-1} = ? \] is, and that is to use the rule that the square of a number times the square root of that number is the number itself.

OpenStudy (phi):

*use the rule that the square *root* of a number times the square root of that number is the number itself.

OpenStudy (anonymous):

4*3 = 12

OpenStudy (phi):

?

OpenStudy (anonymous):

I am lost sorry I am trying to figure out what you are asking

OpenStudy (anonymous):

The square root of -1 is -1 and two negatives equal a positive so I am not sure what you are saying or asking for

OpenStudy (phi):

you said: The square root of -1 is -1 NO! -1*-1 = +1. so -1 is not the root of -1 If we insist that there is a number that is the root of -1, what is it? we say sqrt(-1) is the root of -1, and when we multiply sqrt(-1)*sqrt(-1) we get -1 If this sounds strange, it is. There is no real number that is the root of -1. So people decided to call this strange number i (sqrt(-1)). i*i (by definition)= -1 The very interesting thing is that when you do this, a lot of new math shows up that solves all kinds of problems that could not be solved before. So people got used to using i and imaginary numbers.

OpenStudy (anonymous):

ok so sqrt(-1)*sqrt(-1)= -1

OpenStudy (phi):

yes. and now you can simplify \[ 12\cdot 4\sqrt{3}\cdot \sqrt{-1}\cdot\sqrt{-1} \]

OpenStudy (anonymous):

so will it be 12*4sqrt3*-1 or 12*4sqrt3*i

OpenStudy (phi):

it is the first 12*4*sqrt(3)*-1 but you can do better, right: -1*4*12*sqrt(3) (order can be changed when multiplying) can you finish the multiplication?

OpenStudy (anonymous):

-48sqrt3

OpenStudy (phi):

yes. I hope that how you got the answer makes sense.

OpenStudy (anonymous):

yes so would that be the full answer to my question at the top. because someone else sent x=sqrt-144 = +-12i. and I have no clue how they got that

OpenStudy (phi):

\[ sqrt{-144} = \sqrt{144}\cdot\sqrt{-1}\] now you should REMEMBER \(\sqrt{-1}=i\) and \( \sqrt{144}= ±12 \)

OpenStudy (phi):

*that first thing should be sqrt(-144)

OpenStudy (phi):

they did not multiply by sqrt(-48) , probably because it is not clear in the question

OpenStudy (anonymous):

oic ty soooooo much for being patient and showing me how as well as explaining to me how to do it. Is there anything that I can do for you to show that i am appreciated ,with your help?

OpenStudy (phi):

just keep up the good work! btw, wolfram can do these things, so you can check your answers. for example, see http://www.wolframalpha.com/input/?i=sqrt%28-144%29*sqrt%28-48%29